40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Jun 29 14:13:53 2021, aaron, DailyProgress, Noise Budget, Calibrating PSOMA noise budget 8CD3FEA8-9EE1-4728-8639-A69206A8D098.png
    Reply  Wed Jun 30 14:38:49 2021, aaron, DailyProgress, Noise Budget, Calibrating PSOMA noise budget 210630_PDH_fit.pdf
       Reply  Thu Jul 1 13:39:12 2021, aaron, DailyProgress, Noise Budget, Calibrating PSOMA noise budget 
          Reply  Fri Jul 2 11:29:52 2021, shruti, DailyProgress, Noise Budget, Calibrating PSOMA noise budget 7x
          Reply  Tue Jul 6 14:39:49 2021, shruti, DailyProgress, Noise Budget, Calibrating PSOMA noise budget OLTF_fit.pdfNoise.pdfinitial_calibration_data.zip
             Reply  Thu Jul 8 09:50:41 2021, aaron, DailyProgress, Noise Budget, PSOMA noise budget, does it make sense? Noise.pdfOLTF_fit.pdf
    Reply  Wed Jun 30 17:26:42 2021, aaron, DailyProgress, Noise Budget, Calibrating PSOMA noise budget 6AC149C5-9C0F-42E7-8DE6-AA1F5FB8D9AF.jpeg
       Reply  Fri Jul 30 16:55:38 2021, rana, Electronics, Laser, Delay Line Freq Discriminators 
Message ID: 2775     Entry time: Tue Jul 6 14:39:49 2021     In reply to: 2773     Reply to this: 2776
Author: shruti 
Type: DailyProgress 
Category: Noise Budget 
Subject: Calibrating PSOMA noise budget 

[chris, aaron, shruti]

  • We (Chris, Shruti) noticed that the offset changes on ERC_MON_RATIO when the lights are turned off/ on despite having a long pass filter; it also changes on PDH_CTL_OUT when the slow loop is off. This is probably the main reason why we need to lock the slow controls to an offset of PDH_SET to get the brightest spot.

We (Aaron, Shruti)  re-measured the PDH error signal slope for calibration since the previous measurements were for the settings before the mode-matching was optimized.

drive parameters

A: pk-pk voltage (mV)  B: peak separation time (us) C: sideband crossing separation (us) D: difference in drive at sideband crossings (mV) E: cavity pole (MHz) = 33.59/C*B F: cavity response (mV/MHz) = A/E
1 kHz, 3 Vpp 500 13.6 99.2 480 4.6 109
5 kHz, 3 Vpp 478 2.96 26.1 480 3.8   126

The cavity pole has changed but the peak-peak voltage of the PDH error signal seems roughly the same as measured on Thursday before optimizing the mode-matching. It seems like the different temperature setting we are now at has changed the polarization of light entering the cavity; there is no half-wave plate in the path between the fiber launch and input coupler.

Initial crude noise calibration 

I used the above estimate of the cavity pole and response along with the data measured on Friday to obtain a calibrated noise spectrum (red curve in Attachment 2), then for data above 100 Hz I used the linear estimate of the open loop gain and roll-off shown in Attachment 1 to obtain the blue curve in Attachment 2.

All data and the jupyter notebooks are in Attachment 3

Attachment 1: OLTF_fit.pdf  27 kB  Uploaded Tue Jul 6 18:30:03 2021  | Hide | Hide all
OLTF_fit.pdf
Attachment 2: Noise.pdf  72 kB  Uploaded Tue Jul 6 18:30:15 2021  | Hide | Hide all
Noise.pdf
Attachment 3: initial_calibration_data.zip  339 kB  Uploaded Tue Jul 6 18:31:04 2021
ELOG V3.1.3-